# You Can Wash Hands Better: Accurate Daily Handwashing Assessment with Smartwatches

### v2

@@ Abstract
Hand hygiene is one of the most efficient daily actions to prevent infectious diseases, such as Influenza, Malaria, and skin infections. We have been suggested to wash our hands under professional guidelines to prevent virus infection. However, several surveys show that very few people follow this suggestion. Thus we propose UWash, a wearable solution with smartwatches, to assess handwashing procedures for the purpose of raising users’ awareness and cultivating habits of high-quality handwashing. We address the task of handwashing assessment from readings of motion sensors similar to the action segmentation problem in computer vision, and propose a simple and lightweight two-stream UNet-like network to achieve it effectively. Experiments over 51 subjects show that UWash achieves an accuracy of 92.27% on handwashing gesture recognition, $<$0.5 seconds error on onset/offset detection, and $<$5 points error on gesture scoring in the user-dependent setting, and keeps promising in the user-independent evaluation and the user-independent-location-independent evaluation. UWash even performs well on 10 random passersby in a hospital 9 months later. UWash is the first work that scores the handwashing quality by gesture sequences and is instructive to guide users in promoting hand hygiene in daily life. Code and data are avaliable at https://github.com/aiotgroup/UWash.
@@ 1 Introduction
Back to the conventional approach in the healthcare scenarios (HCS), hospitals and healthcare centers would like to employ auditors to directly observe and assess the handwashing procedures to promote the healthcare workers’ (HCWs) adherence to the guidelines. However, this approach is time-consuming, time-delayed, and costly. For these concerns, many works propose automatic handwashing monitoring systems to promote HCWs’ adherence and awareness [3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Given the above practices, we believe the accessibility of an automatic handwashing monitoring solution in people’s daily life is the next step to promoting ordinary people’s awareness of handwashing.
@@
We propose UWash, an automatic solution with smartwatches, to assess handwashing techniques for the purpose of raising people’s awareness and cultivating habits of high-quality handwashing. As shown in Fig. 1, UWash leverages the inertial measurement unit (IMU) of the smartwatch, i.e., accelerometers and gyroscopes, to detect the start/end time of the handwashing procedures, estimate the duration of each gesture, score the quality of estimated gestures in accordance with WHO guidelines, and finally score the quality of the entire handwashing procedure. With the estimated duration and scores, users could check their handwashing procedures and promote their handwashing techniques accordingly. Compared with existing works, our UWash provides six key features in the usage aspect and functionality aspect that enable it suitable for daily use.
@@
  • With users, no fixed position. For daily use, we enable UWash to assess handwashing procedures along with users, whether they are at home, workplaces, restaurants, etc. In contrast, most of the existing approaches can only work at fixed places where systems are deployed. For example, alcohol sensors [6], pressure sensors [3], ultrasonic hot spots [5], and RFID [4] are embedded into dispensers to count handwashing; cameras [7, 8, 9, 13, 14, 15] and mmWave devices [11] are placed on walls to estimate handwashing.
@@
  • Standalone, no extra sensor. UWash only needs IMU of smartwatches, no extra sensor is required for collaboration. Many existing systems need additional sensors, such as Bluetooth beacons, RFID cards, and Wi-Fi modules on dispensers or on the door, as triggers to detect people’s approaching and start to work [16, 8, 12, 17, 18].
@@
  • Automatic on the start/end detection. UWash can automatically identify the onset and offset of handwashing procedures. In contrast, the official handwashing APP of the Apple Watch and iWash [12] is designed to be awakened by users’ click or voice commands, largely harming the use will and use frequency. AWash [17] leverages the smart faucet or smart foam soap dispenser to detect the onset and offset of handwashing procedures, limiting the assessment places.
@@
  • Elastic for random gesture sequences. Fig. 1 shows 9 gestures of WHO guidelines. Some approaches require the users to rigidly follow the sequences  [4, 16, 19, 12, 13, 14, 20, 21, 15]. However, people always wash their hands in different gesture sequences. UWash is elastic to assessing handwashing procedures no matter the gesture sequences.
  • Details in the duration of each gesture. Many existing works simply count the occurrence of handwashing or report the duration of the entire process [4, 12, 19, 22, 14, 20, 18, 21, 15]. However, the different gestures, shown in Fig. 1, focus on cleaning different parts of the hands. Thus, details of the duration of each gesture are a more important metric for assessing handwashing quality. UWash can report this detailed information.
@@
  • Instructive with scoring. UWash can score the handwashing quality for the purpose of guiding users to improve their handwashing techniques. Since the WHO only provides an overall handwashing recommendation time of 40-60 seconds without specifying the duration for each stage of handwashing, we watched approximately 60 YouTube videos where experts demonstrated handwashing techniques. From these, we carefully selected 12 videos and recorded the duration of each handwashing stage. We then averaged these durations to establish a gold standard for stage durations, which serves as the basis for our scoring, described in Sec. 4.3.
@@
TABLE I: UWash meet all requirements.
@@
We apply the bottom-up schema in UWash for it bypasses segmenting problem in the top-down schema. To tackle the intrinsic false of the bottom-up schema, we no longer conduct gesture recognition over the window, but conduct gesture semantic segmentation over the window. Semantic segmentation in the computer vision community is to align every pixel in an image with an object category, aka pixel-wise classification. In UWash, we adapt U-Net [28], a well-known semantic segmentation network for medical images, to align every sample in a sliding window with a gesture category. As the window slides, every sample in the time-serial IMU readings will have gesture alignment for itself, aka sample-wise gesture classification. To facilitate the gesture semantic segmentation results, Pyramid Pooling Module [29] and channel-wise attention [30] are applied to the U-Net. Fig. 1 shows an example of gesture semantic segmentation on a handwashing procedure. With the estimated procedure segmentation, UWash is able to detect the start and end of handwashing, estimate the duration of each gesture, and score each gesture as well as the entire procedure according to the estimated duration without bells and whistles.
@@
  • Lightweight. UWash models are lightweight, only 496KB.
@@
1. We propose UWash to automatically assess handwashing techniques for people’s daily use, which is the first work that provides all features listed in Table I.
2. We novelly regard the handwashing assessment task as the semantic segmentation task to bypass the intrinsic problem in the bottom-up time-serial classification schema, and design a U-Net variant to achieve it well.
3. We propose a simple approach to obtain a standard of the expected duration of each handwashing gesture following WHO guidelines. With the standard, UWash is the first work that can score the handwashing quality to guide users to improve their handwashing techniques.
4. We collect a dataset with 51 subjects and 5 locations, and conduct an extensive evaluation in settings of user-dependent, cross-user, and cross-location. Besides, we collect a new dataset 9 months later in a hospital to demonstrate the cross-time performance. All datasets and codes have been released for research use.
@@ 2 Related Work
Before the COVID-19 pandemic, hand hygiene has already been crucial to preventing healthcare-associated infections in hospitals. Human mandatory audits are applied to improve the healthcare workers’ compliance with the WHO guidelines. However, this approach is labor-intensive, time-consuming, and costly. Thus automatic handwashing monitoring systems are proposed [3, 4, 5, 6, 7, 8, 9, 10, 11, 12] to facilitate healthcare workers’ adherence. Among them, sensors such as alcohol sensors [6], pressure sensors [3], ultrasonic hotspots [5], and RFID [4] are embedded into the dispensers to simply count the handwashing. RGB cameras [7, 13, 14, 20, 18, 15], depth cameras [8, 9], and mmWave devices [11] are placed on the wall to estimate fine-grained handwashing procedures. Since sensors or devices in these works are required to be deployed next to the handwashing sinks in hospitals, they are not suitable for people’s daily use.
Wearable devices such as wristbands [10, 19], armbands [22], smartwatches [16, 23], and smart rings [21] are also proposed to monitor handwashing in recent years. However, wearing wristbands and armbands in daily life is an extra interaction for users, limiting their widespread use. Considering this, smartwatches are ideal platforms for monitoring handwashing procedures. Unfortunately, current smartwatch-based work [16] cannot detect the start/end time of handwashing effectively, requiring to work along with Bluetooth sensors on dispensers [16] or to be awakened manually [12], where the former limits the use places and the latter harms the use frequency. Lattanzi et al. [23] utilized a single smartwatch to detect hand washing quality. However, this method only supports the monitoring of the overall handwashing duration and cannot measure the duration of individual handwashing steps, limiting the effectiveness of the monitoring. WashRing [21] introduced a hand hygiene monitoring system based on smart ring devices. Although smart rings have the advantages of being compact and unobtrusive to wear, they may interfere with the handwashing process when worn on the fingers. Additionally, the components of hand sanitizers may potentially affect the rings. Besides requirements in user experiences, we believe if the handwashing monitoring system can score every individual handwashing gesture and the entire procedure following WHO guidelines, it will help people to improve their handwashing techniques following the reported scores. Thus we propose UWash to achieve this.
In Table I, we compare UWash with some representative work. The table shows that UWash requires no extra sensor and can work along with users for daily use purposes. In addition, UWash can automatically detect the start/end of the handwashing procedure and can estimate the duration of every gesture regardless of the gesture conducting sequence. Moreover, UWash is the first work with the capability to score the handwashing quality under WHO guidelines.
@@ 3.1 Features of Handwashing
Figure 3: Dilemma. Handwashing patterns emerge in a larger field of view (left). However, a larger field of view may lead to a large prediction error under window-grained prediction when handwashing gestures switch (right). Therefore, we apply the semantic segmentation method for sample-wise prediction in the window to bypass the dilemma.
@@ 4.1 Deep Learning Model
U-Net [28] is a pixel-wise classification network architecture widely used in computer vision tasks of visual semantic segmentation. Later, Temporal U-Net [32] replaces 2-dimensional convolutions in U-Net with 1-dimensional convolutions to conduct sample-wise class alignment on 1-dimensional time-serial data. The U-Net clusters can learn features from data in the wider and wider FoVs with the increase of the convolutional layers (please refer to Section 3.2 for the definition of the FoVs). Further, learned features in large FoVs  (at the high-level layers) and small FoVs (at the low-level layers) are combined with skip connections to promote semantic segmentation performance. Considering these good characters on segmentation, we apply Temporal U-Net with several modifications to achieve our task of the sample-wise handwashing gesture recognition. The network architecture is shown in Fig. 4. Next, we explain the reason why we propose these modifications.
(1) Two-stream. Accelerometers and gyroscopes of smartwatches measure linear accelerations and angular accelerations respectively, describing two physical quantities in different scales. To properly leverage data of the two modalities, we have to do data normalization before merging them for the later task. To conduct automatic modality normalization, we feed raw accelerometer data and raw gyroscope data into the two streams of U-Nets, respectively. We also apply Batch Normalization [33] and Leaky Rectified Linear Unit [34] to facilitate the normalization between the two modalities.
@@
(3) Pyramid Pooling Module (PPM) [29]. PPM is proven to be efficient to harvest features across different FoVs [29]. Therefore, we apply it in the middle of U-Nets, shown in Fig. 4. Before being fed into PPM, feature maps are with the size of $8\times 64$, where 8 and 64 represent temporal dimension and channel dimension, respectively. We use three average pooling operations with window/stride sizes of 8, 4, and 2 on the feature maps, generating three outputs with the size of $1\times 64$, $2\times 64$, and $4\times 64$, respectively. Before concatenating three pooling outputs with the input feature maps along the channel dimension, we use convolution layers with the kernel size of $1\times 1$ to reduce the channel to be of 16, for the channel information balance between the 3 pooling outputs and the input feature maps. Further, we upsample the pooling outputs and concatenate them with the input feature maps, outputting feature maps with the size of $8\times 112$.
@@ 4.3 Handwashing Scoring
We have two empirical assumptions. (1) Since each gesture emphasizes cleaning one part of the hands, we assume each gesture is equally important in handwashing. (2) The quality of cleaning each gesture increases linearly with its duration, and the perfect quality is reached and saturated when the duration is equal to or greater than the professional duration. Given these assumptions, we score handwashing with Equation 2.
$Score=\sum_{i=1}^{9}\frac{100}{9}*min(1,\frac{D_{i}^{e}}{D_{i}^{p}})$ (2)
where $\frac{100}{9}$ is the peak score of each gesture, to match the first assumption; $D_{i}^{e}$ represents the estimated duration of the $i$-th gesture; $min(1,\frac{D_{i}^{e}}{D_{i}^{p}})$ matches the second assumption. Finally, the $Score$ of a handwashing procedure is the sum of estimated scores of all gestures.

### v3

@@ Abstract
Hand hygiene is among the most effective daily practices for preventing infectious diseases such as influenza, malaria, and skin infections. While professional guidelines emphasize proper handwashing to reduce the risk of viral infections, surveys reveal that adherence to these recommendations remains low. To address this gap, we propose UWash, a wearable solution leveraging smartwatches to evaluate handwashing procedures, aiming to raise awareness and cultivate high-quality handwashing habits. We frame the task of handwashing assessment as an action segmentation problem, similar to those in computer vision, and introduce a simple yet efficient two-stream UNet-like network to achieve this goal. Experiments involving 51 subjects demonstrate that UWash achieves 92.27% accuracy in handwashing gesture recognition, an error of $<$0.5 seconds in onset/offset detection, and an error of $<$5 points in gesture scoring under user-dependent settings. The system also performs robustly in user-independent and user-independent-location-independent evaluations. Remarkably, UWash maintains high performance in real-world tests, including evaluations with 10 random passersby at a hospital 9 months later and 10 passersby in an in-the-wild test conducted 2 years later. UWash is the first system to score handwashing quality based on gesture sequences, offering actionable guidance for improving daily hand hygiene. The code and dataset are publicly available at https://github.com/aiotgroup/UWash.
@@ 1 Introduction
In healthcare scenarios, hospitals and healthcare centers would like to employ auditors to directly observe and assess the handwashing procedures to promote the healthcare workers’ (HCWs) adherence to the guidelines. However, this approach is time-consuming, time-delayed, and costly. For these concerns, many works propose automatic handwashing monitoring systems to promote HCWs’ adherence and awareness [3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Given the above practices, we believe the accessibility of an automatic handwashing monitoring solution in people’s daily life is the next step to promoting ordinary people’s awareness of handwashing. Recent advancements in wearable technology have spurred interest in portable handwashing assessment systems. Solutions like WristWash [10] and iWash [12] use wrist-worn devices to evaluate handwashing gestures but rely on predefined motion patterns or manual activation, which reduce usability. Similarly, WashRing [13], employing smart rings, faces issues with potential interference during handwashing. AWash [14] and RFWash [11], on the other hand, utilize infrastructure-based triggers, such as smart faucets or RFID sensors, but lack portability and adaptability across environments.
@@
To address these limitations, We propose UWash, an automatic solution with smartwatches, to assess handwashing techniques for the purpose of raising people’s awareness and cultivating habits of high-quality handwashing. As shown in Fig. 1, UWash leverages the inertial measurement units (IMUs) of the smartwatch, i.e., accelerometers and gyroscopes, to detect the start/end time of the handwashing procedures, estimate the duration of each gesture, score the quality of estimated gestures in accordance with the WHO guidelines, and further score the quality of the entire handwashing procedure. With the estimated duration and scores, users could check their handwashing procedures and promote their handwashing techniques accordingly. Compared with existing works, our UWash provides six key features in the usage aspect and functionality aspect that enable it suitable for daily use.
@@
  • Flexible usage across various locations. UWash is designed to assess handwashing procedures in various environments, including homes, workplaces, and restaurants, offering flexibility for daily use. In contrast, most of the existing approaches can only work at fixed places where systems are deployed. For example, alcohol sensors [6], pressure sensors [3], ultrasonic hot spots [5], and RFID [4] are embedded into dispensers to count handwashing; cameras [7, 8, 9, 15, 16, 17] and mmWave devices [11] are placed on walls to estimate handwashing.
@@
  • Independent operation without additional sensors. UWash relies solely on the IMU of smartwatches, eliminating the need for any extra sensors. In contrast, many existing systems need additional sensors, such as Bluetooth beacons, RFID cards, and Wi-Fi modules on dispensers or on the door, as triggers to detect people’s approaching and start to work [18, 8, 12, 14, 19].
@@
  • Automatic detection of start and end. UWash can automatically identify the onset and offset of handwashing procedures. In contrast, the official handwashing APP of the Apple Watch and iWash [12] require user activation through clicks or voice commands, which can decrease user willingness and frequency of use. AWash [14] leverages the smart faucet or smart foam soap dispenser to detect the onset and offset of handwashing procedures, limiting the assessment places.
@@
  • Flexible for random gesture sequences. Fig. 1 shows 9 gestures of WHO guidelines. Some approaches require the users to strictly follow the sequences  [4, 18, 20, 12, 15, 16, 21, 13, 17]. However, people always wash their hands in different gesture sequences. UWash is adaptable and can assess handwashing procedures regardless of the sequence of gestures used.
  • Detailed duration for each gesture. Most existing methods merely count the number of handwashing occurrences or measure the total duration of the process [4, 12, 20, 22, 16, 21, 19, 17]. However, different gestures target different parts of the hands, making the duration of each specific gesture a crucial metric for assessing handwashing quality. UWash provides this information, ensuring a comprehensive evaluation of handwashing practices.
@@
  • Instructive with scoring. UWash can score the handwashing quality to guide users to improve their handwashing techniques. Since the WHO only provides an overall handwashing recommendation time of 40-60 seconds without specifying the duration for each stage of handwashing, we reviewed approximately 60 YouTube videos featuring expert handwashing demonstrations. From these, we carefully selected 12 videos and recorded the duration of each handwashing stage. We then averaged these durations to establish a gold standard for stage durations, which serves as the basis for our scoring, described in Sec. 4.3.
@@
TABLE I: This table compares different handwashing assessment solutions based on key requirements, such as flexibility, sensor independence, and start/end detection. UWash uniquely meets all these requirements, setting it apart from other methods.
@@
We apply the bottom-up schema in UWash for it bypasses segmenting problem in the top-down schema. To tackle the intrinsic false classification of the bottom-up schema, we no longer conduct gesture recognition over the window, but conduct gesture semantic segmentation over the window. Semantic segmentation in the computer vision community is to align every pixel in an image with an object category, aka pixel-wise classification. In UWash, we adapt U-Net [28], a well-known semantic segmentation network for medical images, to align every sampling point (sample) in a sliding window with a gesture category. As the window slides, every sample in the time-serial IMU readings will have gesture alignment for itself, aka sample-wise gesture classification. To facilitate the gesture semantic segmentation results, Pyramid Pooling Module [29] and channel-wise attention [30] are applied to the U-Net. Fig. 1 shows an example of gesture semantic segmentation on a handwashing procedure. With the estimated procedure segmentation, UWash is able to detect the start and end of handwashing, estimate the duration of each gesture, and score each gesture as well as the entire procedure according to the estimated duration.
@@
  • Lightweight. UWash is lightweight, only 496KB.
@@
(1) We propose UWash to automatically assess handwashing techniques for people’s daily use, which is the first work that provides all features listed in Table I.
(2) We novelly regard the handwashing assessment task as the semantic segmentation task to bypass the intrinsic problem in the bottom-up time-serial classification schema, and design a U-Net variant to achieve it well.
(3) We propose a simple approach to obtain a standard of the expected duration of each handwashing gesture following WHO guidelines. With the standard, UWash is the first work that can score the handwashing quality to guide users to improve their handwashing techniques.
(4) We collect a dataset with 51 subjects and 5 locations, and conduct an extensive evaluation in settings of user-dependent, cross-user, and cross-location. Besides, we collect a new dataset 9 months later in a hospital to demonstrate the cross-time performance. All datasets and codes have been released for research use.
@@ 2 Related Work
Hand hygiene has been crucial to preventing healthcare-associated infections in hospitals. Human mandatory audits are applied to improve the healthcare workers’ compliance with the WHO guidelines. However, this approach is labor-intensive, time-consuming, and costly. Thus automatic handwashing monitoring systems are proposed [3, 4, 5, 6, 7, 8, 9, 10, 11, 12] to facilitate healthcare workers’ adherence. Among them, sensors such as alcohol sensors [6], pressure sensors [3], ultrasonic hotspots [5], and RFID [4] are embedded into the dispensers to simply count the handwashing. RGB cameras [7, 15, 16, 21, 19, 17], depth cameras [8, 9], and mmWave devices [11] are placed on the wall to estimate fine-grained handwashing procedures. Since sensors or devices in these works are required to be deployed next to the handwashing sinks in hospitals, they are not suitable for people’s daily use.
Several wearable and infrastructure-based systems have been proposed for handwashing assessment. WristWash [10] utilizes wrist-worn devices to evaluate handwashing through clip-wise motion recognition but is constrained by predefined sequences, reducing its adaptability to diverse user behaviors. Similarly, iWash [12] relies on smartwatch activation via user interaction, which may discourage frequent use in daily settings.WashRing [13] introduced a hand hygiene monitoring system based on smart ring devices. Although smart rings have the advantages of being compact and unobtrusive to wear, they may interfere with the handwashing process when worn on the fingers.In addition to wearable devices, environmental sensor-based methods such as AWash [14] and RFWash [11] have also been explored. AWash [14] employs smart faucets or soap dispensers to detect handwashing onset and offset, but its reliance on specific infrastructure limits deployment flexibility. RFWash [11] uses RFID tracking for hand hygiene compliance but lacks the granularity to analyze individual gestures or provide detailed feedback. Besides requirements in user experiences, we believe if the handwashing monitoring system can score every individual handwashing gesture and the entire procedure following WHO guidelines, it will help people to improve their handwashing techniques following the reported scores. Thus we propose UWash to achieve this.
In Table I, we compare UWash with some representative work. Unlike infrastructure-based methods like AWash [14] and RFWash [11], which require external sensors or specific setups, UWash relies solely on smartwatch IMU sensors, making it more versatile and accessible for daily use. Compared to wearable-based systems like WristWash [10] and WashRing [13], UWash supports random gesture sequences and avoids the usability challenges posed by predefined patterns or potential interference during handwashing. Moreover, UWash is the first work with the capability to score the handwashing quality, which evaluates the quality of each gesture and the entire procedure based on WHO guidelines. This capability not only ensures compliance but also provides actionable feedback to users, promoting better hygiene practices. Additionally, UWash automates the detection of handwashing start and end events, eliminating the need for manual activation or external triggers. These advantages make UWash a comprehensive and practical solution for hand hygiene monitoring.
@@ 3.1 Features of Handwashing
Since handwashing is a continuous, multi-stage process, where each stage involves repeated actions, while some daily activities may be misclassified as belonging to a particular handwashing phase, the need for specific criteria, such as a period of duration, multiple stages, and repeated actions within each stage, can help filter out these misclassifications.
Figure 3: Handwashing patterns emerge in a larger field of view (left). However, a larger field of view may lead to a large prediction error under window-grained prediction when handwashing gestures switch (right). Therefore, we apply the semantic segmentation method for sample-wise prediction in the window to bypass the dilemma.
@@ 4.1 Deep Learning Model
U-Net [28] is a pixel-wise classification network architecture widely used in computer vision tasks of visual semantic segmentation. Temporal U-Net [32] replaces 2-dimensional convolutions in U-Net with 1-dimensional convolutions to conduct sample-wise class alignment on 1-dimensional time-serial data. The U-Net clusters can learn features from data in the wider FoVs with the increase of the convolutional layers (please refer to Section 3.2 for the definition of the FoVs). Further, learned features in large FoVs  (at the high-level layers) and small FoVs (at the low-level layers) are combined with skip connections to promote semantic segmentation performance. Considering these good characters on segmentation, we apply Temporal U-Net with several modifications to achieve our task of the sample-wise handwashing gesture recognition. The network architecture is shown in Fig. 4. Next, we explain the reason why we propose these modifications.
(1) Two-stream. Accelerometers and gyroscopes of smartwatches measure linear accelerations and angular accelerations respectively, describing two physical quantities in different scales. To properly leverage data of the two modalities, we have to do data normalization before merging them for the later task. The classic Gaussian normalization method requires the mean and the variance of the whole training dataset, which is highly dataset-dependent and sensitive to extremums. To conduct automatic modality normalization, we feed raw accelerometer data and raw gyroscope data into the two streams of U-Nets, respectively. We also apply Batch Normalization [33] and Leaky Rectified Linear Unit [34] to facilitate stable training and fast convergence.
@@
(3) Pyramid Pooling Module (PPM) [29]. PPM is proven to be efficient in harvesting features across different FoVs [29]. Therefore, we apply it in the middle of U-Nets, shown in Fig. 4, to harvest IMUs features of different FoVs in our task. Before being fed into PPM, feature maps are with the size of $8\times 64$, where 8 and 64 represent temporal dimension and channel dimension, respectively. We use three average pooling operations with window/stride sizes of 8, 4, and 2 on the feature maps, generating three outputs with the size of $1\times 64$, $2\times 64$, and $4\times 64$, respectively. Before concatenating three pooling outputs with the input feature maps along the channel dimension, we use convolution layers with the kernel size of $1\times 1$ to reduce the channel to be of 16, for the channel information balance between the 3 pooling outputs and the input feature maps. Further, we upsample the pooling outputs and concatenate them with the input feature maps, outputting feature maps with the size of $8\times 112$.
@@ 4.3 Handwashing Scoring
We have two empirical assumptions. (1) Since each gesture emphasizes cleaning one part of the hands, we assume each gesture is equally important in handwashing. (2) The quality of cleaning of each gesture increases linearly with its duration, and the perfect quality is reached and saturated when the duration is equal to or greater than the professional duration. Given these assumptions, we score handwashing with Equation 2.
$Score=\frac{100}{9}\sum_{i=1}^{9}min(1,\frac{D_{i}^{e}}{D_{i}^{p}})$ (2)
where $\frac{100}{9}$ is the best score of each gesture, to match the first assumption; $D_{i}^{e}$ represents the estimated duration of the $i$-th gesture; $min(1,\frac{D_{i}^{e}}{D_{i}^{p}})$ matches the second assumption. Finally, the $Score$ of a handwashing procedure is the sum of estimated scores of all nine gestures.
@@
Please note that we have taken applying gels into consideration since it is challenging to quantify the scores of various types of sanitization gels (e.g., spray, gel, or foam) and the different ways users apply them (e.g., pressing with fingers or palms).